# Bridged Posterior: Optimization, Profile Likelihood and a New Approach of Generalized Bayes

### v1

@@ 3.2 Diffusion-based Algorithms for Profile Likelihood-based Bridged Posterior
Instead of uniform random walk proposals $G$, informative proposals such as the Metropolis-adjusted Langevin algorithm (MALA) or Hamiltonian Monte Carlo may yield better performance. These algorithms become especially advantageous compared to random walk methods in terms of mixing as the dimension $d$ increases. In this subsection, we first show that gradients (or sub-gradients) are readily available in cases when a profile likelihood is used, and then we discuss its use in the MALA algorithm. Under the bridged posterior, the lack of closed forms for $z$ presents a potential challenge to these methods, leading to intractable gradients or subgradients with respect to $z$. However, for those based on the profile likelihood, this issue can be bypassed entirely. Consider the posterior deriving from (4),
@@
This is due to the envelope theorem. When $z$ may not be differentiable in $\lambda$ but is strictly continuous in $\lambda$, the expression ${\partial g(\zeta,y;\lambda)}/{\partial\lambda}\big{|}_{\zeta=z}$ still holds as a subgradient of $\min_{\zeta}g(\zeta,y;\lambda)$ with respect to $\lambda$ (Rockafellar and Wets, 2009, Theorem 10.49). For completeness, recall a subgradient of $f:\mathbb{R}^{d}\to\mathbb{R}$ at $x\in\mathbb{R}^{d}$ is a vector $v\in\mathbb{R}^{d}$ that satisfies $f(y)\geq f(x)+v^{\top}(y-x)$ for any $y$ in the domain. In subgradient-based MCMC samplers (Tang and Yang, 2022), one typically refers to a local subgradient with inequality held for $y:\|y-x\|\leq\epsilon$ under a sufficiently small $\epsilon>0$. When $f$ is differentiable at $x$, there is a unique subgradient, coinciding with the usual gradient. We use ~$\tilde{\nabla}\log\Pi(\lambda\mid y)$ to denote a subgradient evaluated at point $\lambda$. For reversibility, in the case when there is more than one subgradient at $\lambda$, we impose a constraint that ~$\tilde{\nabla}\log\Pi(\lambda\mid y)$ is chosen as one of the subgradients in a pre-determined way. This constraint is implicitly satisfied in most computing software, for example, most packages will output ~$\tilde{\nabla}|\lambda_{1}|_{1}=0$ when $\lambda_{1}=0$, even though any value $[-1,1]$ is a subgradient. We now describe the MALA algorithm with preconditioning.
@@ 4.1 General Bridged Posterior under Parametric Setting
    The limit $l_{*}$ has positive definite $l_{*}^{\prime\prime}\{\lambda_{0},\hat{z}_{*}(\lambda_{0})\}$ and satisfies $\frac{\partial l_{*}(\lambda_{0},\zeta)}{\partial\zeta}|_{\zeta=\hat{z}_{*}(\lambda_{0})}=0$.
@@
[…]
As mentioned above, the result above shows that fixing $\zeta$ to $z$ does not impact the asymptotic variance of $\lambda$. On the other hand, since $z$ is finite-dimensional and differentiable on $E$, we can use delta method to find out the asymptotic covariance of $z$. For bridged posterior using profile likelihood, we do find lower uncertainty in $\Pi(z\mid y)$ under a bridged posterior compared to $\Pi(\zeta\mid y)$ under an integrated one, as formalized below.
@@ 4.2 Bridged Posterior using Profile Likelihood under Semi-parametric Setting
$P_{\lambda_{0},\zeta_{0}}\biggl{[}\inf_{\|\lambda-\hat{\lambda}_{n}\|\geq\epsilon}\bigl{\{}\hat{l}_{n}(\hat{\lambda}_{n})-\hat{l}_{n}(\lambda)\bigr{\}}\geq\delta\biggr{]}\xrightarrow[n\to\infty]{}1.$ (12)
@@
$\int_{B_{\epsilon}(\lambda_{0})}\pi_{n}(\lambda)\,\textup{d}\lambda\xrightarrow[n\to\infty]{P_{\lambda_{0},\zeta_{0}}}1\text{ for all }\epsilon>0,$ (13)
[…]
4.3 Comparison with Existing BvM Results on Semi-parametric Models
The second category of BvM results relate to canonical Bayesian methodology involving integrated posterior $\Pi(\lambda\mid y)=\int\Pi(\lambda,\textup{d}\zeta\mid y)$ over a non-deterministic $\zeta$. Bickel and Kleijn (2012) proved a BvM result for marginal posterior distribution of $\lambda$ using the LAN property for the marginal likelihood of $\lambda$, which has similar form with (11). On the other hand, they additionally assume that the marginal posterior probability of $\lambda$ inside the neighborhood $B_{M_{n}/\sqrt{n}}(\lambda_{0})$ converges to $1$ for every $M_{n}\to\infty$. This condition is similar to but arguably stronger than (12). Castillo and Rousseau (2015) proved a BvM result on a functional of the parameters, under an essentially necessary no-bias condition which is related to both the likelihood and the prior specification of ($\lambda,\zeta$). To be informative to the interested readers, we provide a detailed summary of their conditions in the supplementary materials.
Before we conclude this section, it is of particular interest to make the connection between our approach (based on partial optimization) and the second category (based on marginalization). It is very encouraging to see the asymptotic equivalence of the two $\sqrt{n}$-adjusted posterior distributions of $\lambda$. It suggests that one can potentially eliminate the computational and interpretability problems created by high-dimensional $\zeta$, by instead transferring to an optimization sub-problem and then applying canonical MCMC on the remaining part for valid uncertainty estimation.
@@ 5.1 Latent Quadratic Exponential Model
We fit the latent quadratic exponential model (6) and the latent normal model (5) to the simulated data. For both models, we assign half-normal $\textup{N}_{+}(0,1)$ prior on $\tau$ and Inverse-Gamma$(2,5)$ prior on $b$. We use random walk Metropolis for the latent quadratic exponential model, and data augmentation Gibbs sampler for the latent normal model (detail provided in the supplementary materials).
We run each MCMC algorithm for $10,000$ iterations and discard the first $2,000$ as burn-ins. The latent quadratic exponential model takes about $8.37$ minutes, and the latent normal model takes about $11.87$ minutes on a $12$-core laptop. Figure 2 compares the mixing of MCMC algorithms for those two models. Clearly, the latent quadratic exponential model mixes better, while taking less runtime. In terms of effective sample size for $(b,\tau)$ per time unit (10 seconds wall time) (ESS/time), the latent quadratic exponential model achieves $0.100$ and $0.129$, while the latent normal model yields only $0.0079$ and $0.0009$.
Refer to caption Traceplot of $b$ Refer to caption Autocorrelation of $b$ Refer to caption Traceplot of $\tau$ Refer to caption Autocorrelation of $\tau$ (a) The MCMC samples produced by the data augmentation Gibbs sampler applied on the latent normal model.
Refer to caption Traceplot of $b$ Refer to caption Autocorrelation of $b$ Refer to caption Traceplot of $\tau$ Refer to caption Autocorrelation of $\tau$ (b) The MCMC samples produced by the random-walk Metropolis applied on the latent quadratic exponential model.
@@
Next, we compare the posterior distributions of parameters $(\tau,b)$. As can be seen in Figure 3, these two distributions show a similar range of $\tau$ and $b$ in the high posterior probability region. Since these two distributions correspond to two distinct models, we do not expect the distributions of $\tau$ or $b$ to match exactly. On the other hand, we can see that the posterior variances are on the same scale, with $\text{Var}(b\mid y)=1.08^{2}$ and $\text{Var}(\tau\mid y)=0.41^{2}$ for the latent normal model, and $\text{Var}(b\mid y)=0.66^{2}$ and $\text{Var}(\tau\mid y)=0.41^{2}$ for the latent quadratic exponential model. We repeat the experiments and compare the variances under different sample sizes. We provide the details in the supplementary materials.
@@ 6 Data Application: Harmonization of Functional Connectivity Graphs
The graph Laplacian $\mathcal{L}^{(s)}=D^{(s)}-A^{(s)}$ is a routinely used one-to-one transform of $A^{(s)}$, where $D^{(s)}$ is a diagonal matrix $D^{(s)}_{ii}=\sum_{j=1}^{R}A_{i,j}^{(s)}$. Compared to the adjacency matrix, the Laplacian enjoys a few appealing properties, namely: (i) $\mathcal{L}^{(s)}$ is always positive semidefinite, (ii) the number of zero eigenvalues equals the number of disjoint component sub-graphs (each known as a community); (iii) the first smallest non-zero eigenvalues quantify the connectivity (normalized graph cut) in each component sub-graph. Because of these properties, we can quantify the difference between two graphs via the geodesic distance in the interior of positive definite cone (Lim et al., 2019). For two positive semidefinite matrices $X$ and $Y$ of equal size, $\text{dist}(X,Y)=\lim_{\eta\to 0_{+}}\Bigl{(}\sum_{j=1}^{R}\log^{2}\bigl{[}\xi_{j}\{(X+I\eta)^{-1}(Y+I\eta)\}\bigr{]}\Bigr{)}^{1/2}$, where $\xi_{j}(\cdot)$ is the $j$-th eigenvalue.
Figure 6(a) plots the pairwise distances between the observed $\mathcal{L}^{(s)}$’s, using three boxplots corresponding to subjects within the diseased group, subjects within the healthy group, and those between the two groups. Though we see that the within-group distances have slightly smaller means than the between-group distances, there is significant overlap among the three boxplots. This is understandable since each observed $\mathcal{L}^{(s)}$ is of rank $(R-1)$ corresponding to having no disjoint components (equivalent to one community only). It is reasonable to take a reduced-rank smoothing by solving the following convex problem:
$Z^{(s)}={\arg\min}_{\zeta}\ \frac{1}{2}\|\mathcal{L}^{(s)}-\zeta\|^{2}+\tilde{\lambda}_{s}\|\zeta\|_{*}$
@@
where $\|\zeta\|_{*}$ is the nuclear norm of matrix $\zeta$ (the sum of singular values of $\zeta$), and $\tilde{\lambda}_{s}>0$ is the tuning parameter. As the result, each $Z^{(s)}$ with rank $(R-K^{(s)})$ is the graph Laplacian of a $K^{(s)}$-community graph, where $K^{(s)}$ is monotonically non-decreasing in $\tilde{\lambda}_{s}>0$.
Note that treating $Z^{(s)}$ here as a latent variable is especially appealing—if we had viewed it as a parameter in $\mathbb{R}^{n\times n}$, we would incur a high modeling and computational cost. The question boils down to a meaningful choice of $\tilde{\lambda}_{s}$. Due to the heterogeneity of $\mathcal{L}^{(s)}$, the same value of $\tilde{\lambda}_{s}=\tilde{\lambda}_{s^{\prime}}$ may yield quite different $Z^{(s)}$ and $Z^{(s^{\prime})}$. As a result, for the purpose of data harmonization, instead of assigning an independent prior or equal value on $\tilde{\lambda}_{s}$, we assign a dependent likelihood based on the pairwise distances among $Z^{(s)}$:
@@
To calculate $Z^{(s)}$ at different $\tilde{\lambda}_{s}$, we use the alternating direction method of multipliers (ADMM) algorithm (details are provided in the supplementary materials). To facilitate computation, for each $\tilde{\lambda}_{s}$, we assign a discrete uniform equally spread over $10$ values in $(0,5]$, so that the possible values of $Z^{(s)}$ as well as their associated pairwise geodesic distances can be precomputed before running MCMC. We specify an Inverse-Gamma$(2,1)$ prior for $\sigma^{2}$ and Inverse-Gamma$(2,1)$ prior for $\tau$. Running MCMC for $10,000$ iterations takes 46.5 minutes on a 12-core laptop; the first $2,000$ samples are discarded as burn-in.
@@
Refer to caption Healthy subject 1. Refer to caption Healthy subject 2. Refer to caption Diseased subject 1. Refer to caption Diseased subject 2. (a) Smoothed adjacency matrices for four subjects.
Refer to caption Top-left $64\times 64$: healthy subjects, bottom-right $102\times 102$: diseased subjects. (b) Posterior mean of the pairwise geodesic distance among the smoothed Laplacians.

### v2

@@ 3.2 Diffusion-based Algorithms for Profile Likelihood-based Bridged Posterior
Instead of uniform random walk proposals $G$, informative proposals such as the Metropolis-adjusted Langevin algorithm (MALA) or Hamiltonian Monte Carlo may yield better performance. These algorithms become especially advantageous compared to random walk methods in terms of mixing as the dimension $d$ increases.
In this subsection, we first show that gradients (or sub-gradients) are readily available in cases when a profile likelihood is used, and then we discuss its use in the MALA algorithm. Under the bridged posterior, the lack of closed forms for $z$ presents a potential challenge to these methods, leading to intractable gradients or subgradients with respect to $z$. However, for those based on the profile likelihood, this issue can be bypassed entirely. Consider the posterior deriving from (5),
@@
This is due to the envelope theorem.
When $z$ may not be differentiable in $\lambda$ but is strictly continuous in $\lambda$, the expression ${\partial g(\zeta,y;\lambda)}/{\partial\lambda}\big{|}_{\zeta=z}$ still holds as a subgradient of $\min_{\zeta}g(\zeta,y;\lambda)$ with respect to $\lambda$ (Rockafellar and Wets, 2009, Theorem 10.49). For completeness, recall a subgradient of $f:\mathbb{R}^{d}\to\mathbb{R}$ at $x\in\mathbb{R}^{d}$ is a vector $v\in\mathbb{R}^{d}$ that satisfies $f(y)\geq f(x)+v^{\top}(y-x)$ for any $y$ in the domain. In subgradient-based MCMC samplers (Tang and Yang, 2024), one typically refers to a local subgradient with inequality held for $y:\|y-x\|\leq\epsilon$ under a sufficiently small $\epsilon>0$. When $f$ is differentiable at $x$, there is a unique subgradient, coinciding with the usual gradient.
We use ~$\tilde{\nabla}\log\Pi(\lambda\mid y)$ to denote a subgradient evaluated at point $\lambda$. For reversibility, in the case when there is more than one subgradient at $\lambda$, we impose a constraint that ~$\tilde{\nabla}\log\Pi(\lambda\mid y)$ is chosen as one of the subgradients in a pre-determined way. This constraint is implicitly satisfied in most computing software, for example, most packages will output ~$\tilde{\nabla}|\lambda_{1}|_{1}=0$ when $\lambda_{1}=0$, even though any value $[-1,1]$ is a subgradient. We now describe the MALA algorithm with preconditioning.
@@ 4.1 General Bridged Posterior under Parametric Setting
    The limit $l_{*}$ has positive definite $-l_{*}^{\prime\prime}\{\lambda_{0},\hat{z}_{*}(\lambda_{0})\}$ and satisfies $\frac{\partial l_{*}(\lambda_{0},\zeta)}{\partial\zeta}|_{\zeta=\hat{z}_{*}(\lambda_{0})}=0$.
@@
Under (A1A4), there is a sequence $\lambda_{n}\to\lambda_{0}$ such that $\hat{l}_{n}^{\prime}(\lambda_{n})=0$ for all $n$ large enough, $\hat{l}_{n}(\lambda_{n})\to\hat{l}_{*}(\lambda_{0})$ where $\hat{l}_{*}(\lambda)=l_{*}\{\lambda,\hat{z}_{*}(\lambda)\}$. Further, letting $q_{n}$ be the density of $\sqrt{n}(\lambda-\lambda_{n})$ when $\lambda\sim\Pi_{n}(\lambda\mid y)$, and $\mathcal{N}$ the normal density, we have the total variational distance $\text{d}_{\text{TV}}\bigl{\{}q_{n},\mathcal{N}\bigl{(}0,H_{0}^{-1}\bigr{)}\bigr{\}}\xrightarrow[n\to\infty]{a.s.[y_{1:n}]}0$ with $H_{0}=\hat{l}_{*}^{\prime\prime}(\lambda_{0})$.
The result above shows that fixing $\zeta$ to $z$ does not impact the asymptotic variance of $\lambda$. On the other hand, since $z$ is finite-dimensional and differentiable on $E$, we can use the delta method to find out the asymptotic covariance of $z$. For bridged posterior using profile likelihood, we do find lower uncertainty in $\Pi(z\mid y)$ under a bridged posterior compared to $\Pi(\zeta\mid y)$ under an integrated one, as formalized below.
@@
Remark 4.
In the proof of Corollary 1, we show that for the bridged posterior based on profile likelihood, the inverse asymptotic variance $H_{0}=-\hat{l}^{\prime\prime}_{*}(\lambda_{0})=-l_{*,\lambda_{0}\lambda_{0}}+l_{*,\lambda_{0}\zeta_{0}}l_{*,\zeta_{0}\zeta_{0}}^{-1}l_{*,\zeta_{0}\lambda_{0}}$, where $l_{*,\lambda_{0}\lambda_{0}},l_{*,\zeta_{0}\zeta_{0}},l_{*,\zeta_{0}\lambda_{0}},l_{*,\lambda_{0}\zeta_{0}}$ are the second partial derivatives of $l_{*}$ evaluated at $\lambda=\lambda_{0},\zeta=\zeta_{0}=\hat{z}_{*}(\lambda_{0})$. On the other hand, for the integrated posterior (marginal posterior), letting $\tilde{Q}_{n}$ denote the posterior distribution of $\sqrt{n}(\lambda-\lambda_{n})$ when $\lambda\sim\tilde{\Pi}(\lambda\mid y)\propto\pi_{0}(\lambda)\int L(y,\textup{d}\zeta;\lambda)$, the BvM theorem (Miller, 2021) states that $\text{d}_{\text{TV}}\bigl{\{}\tilde{Q}_{n},\mathcal{N}\bigl{(}0,\tilde{H}_{0}^{-1}\bigr{)}\bigr{\}}\xrightarrow[n\to\infty]{a.s.[y_{1:n}]}0$ where $\tilde{H}_{0}^{-1}$ is the $\lambda$-block of the inverse full Fisher information $-\bigl{[}{l}_{*}^{\prime\prime}(\lambda_{0},\zeta_{0})\bigr{]}^{-1}$. Block matrix inversion then shows that $H_{0}=\tilde{H}_{0}$.
@@ 4.2 Bridged Posterior using Profile Likelihood under Semi-parametric Setting
$P_{\lambda_{0},\zeta_{0}}\biggl{[}\inf_{\|\lambda-\hat{\lambda}_{n}\|\geq\epsilon}\bigl{\{}\hat{l}_{n}(\hat{\lambda}_{n})-\hat{l}_{n}(\lambda)\bigr{\}}\geq\delta\biggr{]}\xrightarrow[n\to\infty]{}1.$
@@
$\int_{B_{\epsilon}(\lambda_{0})}\pi_{n}(\lambda)\,\textup{d}\lambda\xrightarrow[n\to\infty]{P_{\lambda_{0},\zeta_{0}}}1\text{ for all }\epsilon>0,$
and letting $q_{n}$ be the density of $\sqrt{n}(\lambda-\hat{\lambda}_{n})$, we have $\text{d}_{\text{TV}}\bigl{\{}q_{n},\mathcal{N}\bigl{(}0,H_{0}^{-1}\bigr{)}\bigr{\}}\xrightarrow[n\to\infty]{P_{\lambda_{0},\zeta_{0}}}0$.
Remark 5.
We provide a detailed comparison with existing BvM results on semi-parametric models in Appendix C. Here, we compare the asymptotic variance of the bridged posterior with that of the integrated posterior. According to the BvM theorem for the integrated posterior (Bickel and Kleijn, 2012, Theorem 2.1), under certain regularity conditions, letting $\tilde{Q}_{n}$ denote the posterior probability distribution of $\sqrt{n}(\lambda-\lambda_{0})$ where $\lambda\sim\tilde{\Pi}(\lambda\mid y)\propto\pi_{0}(\lambda)\int L(y,\textup{d}\zeta;\lambda)$, we have $\text{d}_{\text{TV}}\bigl{\{}\tilde{Q}_{n},\mathcal{N}\bigl{(}\tilde{\Delta}_{n},\tilde{I}_{0}^{-1}\bigr{)}\bigr{\}}\xrightarrow[n\to\infty]{P_{\lambda_{0},\zeta_{0}}}0$, where $\tilde{\Delta}_{n}=\sqrt{n}\tilde{I}_{0}^{-1}\mathscr{Q}\dot{l}_{n}(\lambda_{0},\zeta_{0})$. Here, $\tilde{I}_{0}$ and $\mathscr{Q}\dot{l}_{n}(\lambda_{0},\zeta_{0})$ denote the efficient Fisher information and the efficient score function, respectively, as defined earlier. A key assumption in their theory is the existence of a least favorable model, which ensures that the submodel satisfies $\partial\tilde{l}_{n}(t,\lambda_{0},\zeta_{0})/{\partial t}|_{t=\lambda_{0}}=\mathscr{Q}\dot{l}_{n}(\lambda_{0},\zeta_{0}),$ and under this condition, the efficient Fisher information is given by $\tilde{I}_{0}=-\mathbb{E}_{\lambda_{0},\zeta_{0}}\bigl{\{}\partial^{2}\tilde{l}_{n}(t,\lambda_{0},\zeta_{0})/{\partial t^{2}}|_{t=\lambda_{0}}\bigr{\}}$. This formulation coincides with our assumption in (B1), under which we also have inverse asymptotic variance $H_{0}=\tilde{I}_{0}$.
@@ 5.1 Latent Quadratic Exponential Model
We fit the latent quadratic exponential model (LABEL:eq:latent_quadratic) and the latent normal model (6) to the simulated data. For both models, we assign half-normal $\textup{N}_{+}(0,1)$ prior on $\tau$ and Inverse-Gamma$(2,5)$ prior on $b$. We use random walk Metropolis for the latent quadratic exponential model, and data augmentation Gibbs sampler for the latent normal model (detail provided in Appendix D).
We run each MCMC algorithm for $10,000$ iterations and discard the first $2,000$ as burn-ins. The latent quadratic exponential model takes about $8.4$ minutes, and the latent normal model takes about $11.9$ minutes on a $12$-core processor. Figure 2 compares the mixing of MCMC algorithms for those two models. Clearly, the latent quadratic exponential model mixes better, while taking less runtime. In terms of effective sample size for $(b,\tau)$ per time unit (10 seconds wall time) (ESS/time), the latent quadratic exponential model achieves $0.059$ for $b$ and $0.106$ for $tau$, while the latent normal model yields only $0.0087$ and $0.0008$, respectively.
Traceplot of $b$
Autocorrelation of $b$
Traceplot of $\tau$
Autocorrelation of $\tau$
(a) The MCMC samples produced by the data augmentation Gibbs sampler applied on the latent normal model.
Traceplot of $b$
Autocorrelation of $b$
Traceplot of $\tau$
Autocorrelation of $\tau$
(b) The MCMC samples produced by the random-walk Metropolis applied on the latent quadratic exponential model.
@@
Next, we compare the posterior distributions of parameters $(\tau,b)$. As can be seen in Figure 3, these two distributions show similar ranges of $\tau$ and $b$ in the high posterior probability region. Since these two distributions correspond to two distinct models, we do not expect the distributions of $\tau$ or $b$ to match exactly. On the other hand, we can see that the posterior variances are on the same scale, with $\text{Var}(b\mid y)=0.60^{2}$ and $\text{Var}(\tau\mid y)=0.65^{2}$ for the latent normal model, and $\text{Var}(b\mid y)=0.94^{2}$ and $\text{Var}(\tau\mid y)=0.51^{2}$ for the latent quadratic exponential model. We repeat the experiments and compare the variances under different sample sizes. Additionally, we compare these empirically to two posterior approximation algorithms, integrated nested Laplace approximation and mean-field variational inference, with details in Section E.1 and Section E.2.
@@ 6 Data Application: Harmonization of Functional Connectivity Graphs
The graph Laplacian $\mathcal{L}^{(s)}=D^{(s)}-A^{(s)}$ is a routinely used one-to-one transform of $A^{(s)}$, where $D^{(s)}$ is a diagonal matrix $D^{(s)}_{ii}=\sum_{j=1}^{R}A_{i,j}^{(s)}$. Compared to the adjacency matrix, the Laplacian enjoys a few appealing properties, namely: (i) $\mathcal{L}^{(s)}$ is always positive semidefinite, (ii) the number of zero eigenvalues equals the number of disjoint component sub-graphs (each known as a community); (iii) the smallest non-zero eigenvalues quantify the connectivity (normalized graph cut) in each component sub-graph. Because of these properties, we can quantify the difference between two graphs via the geodesic distance in the interior of the positive definite cone (Lim et al., 2019). For two positive semidefinite matrices $X$ and $Y$ of equal size, $\text{dist}(X,Y)=\lim_{\eta\to 0_{+}}\Bigl{(}\sum_{j=1}^{R}\log^{2}\bigl{[}\xi_{j}\{(X+I\eta)^{-1}(Y+I\eta)\}\bigr{]}\Bigr{)}^{1/2}$, where $\xi_{j}(\cdot)$ is the $j$-th eigenvalue.
Figure 6(a) plots the pairwise distances between the observed $\mathcal{L}^{(s)}$ using three boxplots corresponding to subjects within the diseased group, subjects within the healthy group, and those between the two groups. While we see that the within-group distances have slightly smaller means than the between-group distances, there is significant overlap among the three boxplots. This is understandable since each observed $\mathcal{L}^{(s)}$ is of rank $(R-1)$, corresponding to having no disjoint components (i.e. only one community), motivating a reduced-rank modeling approach.
We therefore consider each $\mathcal{L}^{(s)}$ as generated near a manifold $\mathcal{M}^{(s)}$, each with a Gaussian-type density proportional to $\exp\left\{-{\text{dist}(\mathcal{L}^{(s)},\mathcal{M}^{(s)})^{2}}/{2\sigma^{2}}\right\}$. This manifold $\mathcal{M}^{(s)}$ is given by the intersection between the space of Laplacians and the nuclear norm ball of radius $r^{(s)}>0$:
$\mathcal{M}^{(s)}=\biggl{\{}\|\zeta\|_{*}\leq r^{(s)},\zeta\in\mathbb{R}^{R\times R}\mid\zeta_{i,i}=-\sum_{j:j\neq i}\zeta_{i,j},\zeta_{i,j}=\zeta_{j,i}\leq 0\text{ for }i\neq j\biggr{\}}.$
Here we denote $\|\zeta\|_{*}$ the nuclear norm of matrix $\zeta$, given by the sum of singular values of $\zeta$. Akin to how the $\ell_{1}$ ball promotes sparsity, the boundary of the nuclear norm ball coincides with matrices of low rank. Since we do not know $r^{(s)}$ (hence do not know $\mathcal{M}^{(s)}$ completely), we want to quantify the uncertainty in the bridged posterior framework.
The distance to $\mathcal{M}^{(s)}$ can be computed by first solving for the projection:
$Z^{(s)}={\arg\min}_{\zeta}\ \frac{1}{2}\|\mathcal{L}^{(s)}-\zeta\|_{F}^{2}+\tilde{\lambda}_{s}(\|\zeta\|_{*}-r^{(s)})$ (12)
@@
where $\tilde{\lambda}_{s}\geq 0$ is a Lagrange multiplier, and $\|Z^{(s)}\|_{*}\leq r^{(s)}$; then we have $\text{dist}(\mathcal{L}^{(s)},\mathcal{M}^{(s)})=\|\mathcal{L}^{(s)}-Z^{(s)}\|_{F}$. Although we do not know $r^{(s)}$ in advance, we know if we were given the value of $\tilde{\lambda}_{s}>0$, then $\tilde{Z}^{(s)}={\arg\min}_{\zeta}\ \frac{1}{2}\|\mathcal{L}^{(s)}-\zeta\|_{F}^{2}+\tilde{\lambda}_{s}\|\zeta\|_{*}$ would be the same solution of (12) given $r^{(s)}=\|\tilde{Z}^{(s)}\|_{*}$ (provided $\mathcal{L}^{(s)}$ is outside $\mathcal{M}^{(s)}$). Therefore, assigning a prior to $\tilde{\lambda}_{s}>0$ is equivalent to assigning a prior on $r^{(s)}$.
The question boils down to how to meaningfully model the collection of $\tilde{\lambda}_{s}$.
Due to the heterogeneity of $\mathcal{L}^{(s)}$, the same value of $\tilde{\lambda}_{s}=\tilde{\lambda}_{s^{\prime}}$ may yield quite different $Z^{(s)}$ and $Z^{(s^{\prime})}$. As a result, for the purpose of data harmonization, instead of assigning independent priors or setting equal values for $\tilde{\lambda}_{s}$, we assign a dependent likelihood based on the pairwise distances among $Z^{(s)}$:
@@
To calculate $Z^{(s)}$ at different $\tilde{\lambda}_{s}$, we use the alternating direction method of multipliers (ADMM) algorithm (details are provided in Appendix F). To facilitate computation, for each $\tilde{\lambda}_{s}$, we assign a discrete uniform equally spread over $10$ values in $(0,5]$, so that the possible values of $Z^{(s)}$ as well as their associated pairwise geodesic distances can be precomputed before running MCMC. We specify an Inverse-Gamma$(2,1)$ prior for $\sigma^{2}$ and Inverse-Gamma$(2,1)$ prior for $\tau$. Running MCMC for $10,000$ iterations takes 46.5 minutes on a 12-core laptop; the first $2,000$ samples are discarded as burn-in.
@@
Healthy subject 1.
Healthy subject 2.
Diseased subject 1.
Diseased subject 2.
(a) Smoothed adjacency matrices for four subjects.
Top-left $64\times 64$: healthy subjects, bottom-right $102\times 102$: diseased subjects.
(b) Posterior mean of the pairwise geodesic distance among the smoothed Laplacians.
@@
Remark 6.
Using bridged posterior has a clear computational advantage over the canonical model with the smoothed matrices being completely random. It is conceivably difficult to sample from the canonical posterior, as it would involve drawing $S$ many high-dimensional matrices. From a data harmonization perspective, practitioners typically want to produce a single matrix per subject; hence, the bridged posterior (using projection) provides a straightforward solution.